Q. Construct the truth tables for (p & q here are not above statement)

a) p V ~q
b) ~p V ~q
c) q -->p
d) ~(p-->q)-->r
1
Expert's answer
2020-12-11T06:10:11-0500

"\begin{matrix}
 P & q&\tilde q &pv\tilde q\\
 T&T&F&T\\
 T&F&T&T\\
 F&T&F&F\\
 F&F&T&T\\
\end{matrix}\\

\begin{matrix}
 P & \tilde p& q&\tilde q &\tilde pv\tilde q\\
 T&F&T&F&F\\
 T&F&F&T&T\\
 F&T&T&F&T\\
 F&T&F&T&T\\
\end{matrix}\\

\begin{matrix}
 P & q&q\implies p\\
 T&T&T\\
 T&F&T\\
 F&T&F\\
 F&F&T\\
\end{matrix}\\

\begin{matrix}
 P & q&r&p\implies q&\tilde( p \implies q)&\tilde(p\implies q)\implies r\\
 T&T&T&T&F&T\\
 T&T&F&T&F&T\\
 T&F&T&F&T&T\\
 T&F&F&F&T&F\\
 F&T&T&T&F&T\\
 F&T&F&T&F&T\\
 F&F&T&T&F&T\\
 F&F&F&T&F&T\\
\end{matrix}\\"